# Grounding Multimodal LLMs to Embodied Agents that Ask for Help with Reinforcement Learning

### v3

@@ 3 Ask-to-Act task
Our goal is to build embodied agents capable of completing under-specified tasks by interacting with users to ask clarification questions to resolve ambiguity and identify preferences of a user in partially observable settings. To study building such agents, we propose the Ask-to-Act task as shown in Fig.˜1, where an agent is spawned randomly in an unseen indoor environment and tasked with fetching single or multiple instances of objects and placing them on a designated receptacle using an under-specified language instruction. Consider the example shown in Fig.˜1: the user wants a specific green cup but, instead of describing it in detail, asks the agent, “Bring the cup and place it on coffee table”. While searching, the agent finds multiple cups on the kitchen counter and the table - two red cups, a white cup, and two green cups. Since the user’s intent is unclear, an optimal agent must ask the fewest clarification questions to disambiguate the requested object. To study an agent’s ability to reason about ambiguity and preferences of the user for such under-specified tasks, we synthetically create tasks of varying difficulty that require identifying user preferences or resolving different types of ambiguities, e.g. about variations in object appearance, location, placement location, size or composition of these axes. Fig.˜1 and Fig.˜5 shows examples of tasks in our dataset, and below we describe each task and evaluation criteria in detail:
@@
3.) Spatial Reasoning: Agent must differentiate between objects with similar appearance based on spatial location to identify target object (e.g., red cup on the light table, or kitchen counter, or dark table as shown in Fig.˜5 (b) in Appendix˜A).
4.) Object Size: The agent must distinguish between objects of different sizes but similar appearance and location to identify target object (e.g., a large red bowl vs. a small red bowl, shown in Fig.˜5 (e) in Appendix˜A).
5.) Compositional Ambiguity Reasoning: A combination of attribute recognition, object size, and spatial reasoning, where the agent must consider appearance, geometry and location to ask minimum number of clarification questions to identify the target object (as shown by example in Fig.˜5 (c, f) in Appendix˜A).
@@ Appendix A Dataset
In Fig.˜5, we show additional examples of different types of ambiguities in our Ask-to-Act dataset. Specifically we show 3 additional examples for:
  • Attribute Recognition: The agent must reason about appearance-based attributes such as color and object category to disambiguate the object user is looking for (e.g., a red bowl and white bowl as shown in Fig.˜5 (a)).
  • Object Size: The agent must distinguish between objects of different sizes but similar appearance and location to disambiguate the object user is looking for (e.g., a large red bowl vs. a small red bowl as shown in Fig.˜5 (b)) .
  • Attribute/Spatial Reasoning and Object Size: A combination of attribute and spatial reasoning with object size, where the agent must account for both appearance or location and size to effectively disambiguate the target object the user is looking for. For the example shown in Fig.˜5 (c), the agent needs to reason about different colored shoes (blue, white, and orange shoe) which are located on the blue cabinet and the sofa and some of those have different size e.g. large blue shoe on blue cabinet vs. small blue shoe on sofa and large white shoe on sofa vs. small white shoe on blue cabinet.
@@
Figure 5: Dataset. Examples of different types of synthetically created ambiguities in Ask-to-Act tasks.

### v5

@@ 3 Ask-to-Act task
Our goal is to build embodied agents capable of completing under-specified tasks by interacting with users to ask clarification questions to resolve ambiguity and identify preferences of a user in partially observable settings. To study building such agents, we propose the Ask-to-Act task as shown in Fig.˜1, where an agent is spawned randomly in an unseen indoor environment and tasked with fetching single or multiple instances of objects and placing them on a designated receptacle using an under-specified language instruction. Consider the example shown in Fig.˜1: the user wants a specific green cup but, instead of describing it in detail, asks the agent, “Bring the cup and place it on coffee table”. While searching, the agent finds multiple cups on the kitchen counter and the table - two red cups, a white cup, and two green cups. Since the user’s intent is unclear, an optimal agent must ask the fewest clarification questions to disambiguate the requested object. To study an agent’s ability to reason about ambiguity and preferences of the user for such under-specified tasks, we synthetically create tasks of varying difficulty that require identifying user preferences or resolving different types of ambiguities, e.g. about variations in object appearance, location, placement location, size or composition of these axes. Fig.˜1 and Fig.˜5(a) shows examples of tasks in our dataset, and below we describe each task and evaluation criteria in detail:
@@
3.) Spatial Reasoning: Agent must differentiate between objects with similar appearance based on spatial location to identify target object (e.g., red cup on the light table, or kitchen counter, or dark table as shown in Fig.˜5(a) (b) in Appendix˜A).
4.) Object Size: The agent must distinguish between objects of different sizes but similar appearance and location to identify target object (e.g., a large red bowl vs. a small red bowl, shown in Fig.˜5(a) (e) in Appendix˜A).
5.) Compositional Ambiguity Reasoning: A combination of attribute recognition, object size, and spatial reasoning, where the agent must consider appearance, geometry and location to ask minimum number of clarification questions to identify the target object (as shown by example in Fig.˜5(a) (c, f) in Appendix˜A).
@@ Appendix A Dataset
In Fig.˜5(a), we show additional examples of different types of ambiguities in our Ask-to-Act dataset. Specifically we show 3 additional examples for:
  • Attribute Recognition: The agent must reason about appearance-based attributes such as color and object category to disambiguate the object user is looking for (e.g., a red bowl and white bowl as shown in Fig.˜5(a) (a)).
  • Object Size: The agent must distinguish between objects of different sizes but similar appearance and location to disambiguate the object user is looking for (e.g., a large red bowl vs. a small red bowl as shown in Fig.˜5(a) (b)) .
  • Attribute/Spatial Reasoning and Object Size: A combination of attribute and spatial reasoning with object size, where the agent must account for both appearance or location and size to effectively disambiguate the target object the user is looking for. For the example shown in Fig.˜5(a) (c), the agent needs to reason about different colored shoes (blue, white, and orange shoe) which are located on the blue cabinet and the sofa and some of those have different size e.g. large blue shoe on blue cabinet vs. small blue shoe on sofa and large white shoe on sofa vs. small white shoe on blue cabinet.
@@
(a) Dataset. Examples of different types of synthetically created ambiguities in Ask-to-Act tasks.